If Tan{(π/4)+x}+tan{(π/4)-x}=aProve that Tan^3{(π/4)+x}+tan^3{(π/4)-x}=a^3-3a

Grade:11

1 Answers

Sayam
6 years ago
Tan{(π/4)+x}+tan{(π/4)-x}=a
now, Tan^3{(π/4)+x}+tan^3{(π/4)-x}=
*Tan{(π/4)+x} by applying a3+b3=(a+b)(a2-ab+b2)Tan{(π/4)+x}Tan2{(π/4)-x}-Tan2{(π/4)+x}+[Tan{(π/4)+x}+tan{(π/4)-x}][
=a[Tan2{(π/4)+x}+Tan2{(π/4)-x}-1]    
=a[[Tan{(π/4)+x}+tan{(π/4)-x}]2-3]
=a[a2-3]=a3-3a